3 bags and 4 pens together cost rs.257 whereas 4 bags and 3 pens together cost rs.324 .find the total cost of 1 bag and 10 pens.

Answer by checkley79(3341)

3B+4P=257 MULTIPLY BY 3.
4B+3P=324 MULTIPLY BY -4 & ADD.
9B+12P=771
-16B-12P=-1296
--------------------------
-7B=-525
B=-525/-7
B=75 rs. COST OF 1 BAG.
3*75+4P=257
225+4P=257
4P=257-225
4P=32
P=32/4
P=8 rs.
1B+10P
1*75+10*8=75+80=155 rs. ANS.